2.00 × 1020 electrons flow through a cross section of a 3.50-mm-diameter iron wire in 6.00 s .

Part A

What is the electron drift speed?

___________ μm/s  
Answer #1

The drift velocity is

Vd = I / nqA

Vd = I / nq*pir^2

Vd = Q/t / nq*pir^2

Vd = [(2 x 10^20 electron x 1.60 x 10^-19 C/1 electron) / 6 sec] / [(8.5 x 10^28 m^3 x 1.6 x 10^-19 x pi x (1.75 x 10^-3)^2

Vd = 5.333 / 130847.33

Vd = 0.00004076 m

Vd = 40.76 um